OpenStudy (anonymous):

Simplify the expression. 9xy^2 – 11xy^2

OpenStudy (anonymous):

9x9=81 11x11=121

OpenStudy (anonymous):

how did you come up with that @009infinity

OpenStudy (anonymous):

you have in both terms xy^2 so you just subtract the coefficients

OpenStudy (anonymous):

geesh these confuss me lol

OpenStudy (anonymous):

another example 3x-5x=(3-5)x=-2x

OpenStudy (anonymous):

ok ty i think i might have it not promised though i am a old geiser lol

OpenStudy (anonymous):

its never too late,:)